### v1

# Graph-Dependent Regret Bounds in Multi-Armed Bandits with Interference

Abstract
Multi-armed bandits (MABs) are frequently used for online sequential decision-making in applications ranging from recommending personalized content to assigning treatments to patients. A recurring challenge in the applicability of the classic MAB framework to real-world settings is ignoring interference, where a unit’s outcome depends on treatment assigned to others. This leads to an exponentially growing action space, rendering standard approaches computationally impractical. We study the MAB problem under network interference, where each unit’s reward depends on its own treatment and those of its neighbors in a given interference graph. We propose a novel algorithm that uses the local structure of the interference graph to minimize regret. We derive a graph-dependent upper bound on cumulative regret showing that it improves over prior work. Additionally, we provide the first lower bounds for bandits with arbitrary network interference, where each bound involves a distinct structural property of the interference graph. These bounds demonstrate that when the graph is either dense or sparse, our algorithm is nearly optimal, with upper and lower bounds that match up to logarithmic factors. We complement our theoretical results with numerical experiments, which show that our approach outperforms baseline methods.
Machine Learning, ICML
1 Introduction
Multi-armed bandits (MABs) have become a fundamental tool for online decision-making in a variety of applications (Bouneffouf et al., 2020; Tewari & Murphy, 2017; Lattimore & Szepesvári, 2020; Agarwal et al., 2024; Bubeck et al., 2009; Shahverdikondori et al., 2025). Although classic MABs perform very well in certain applications, they are susceptible to systematic errors that lead to suboptimal results when interference is present among units. Interference occurs when the outcome (reward) of one unit depends not only on its treatment but also on the treatments of others. For instance, the vaccination status of an individual can affect the likelihood that others will fall ill, or a personal advertisement shown to an individual may get shared with their friends, causing them to purchase the item.
Such interference poses significant challenges to traditional experimental design and sequential learning frameworks, which often assume independent unit responses, commonly referred to as the Stable Unit Treatment Value Assumption (SUTVA) (Rubin, 1980). Breakdown of SUTVA in systems with cross-unit dependencies complicates the applicability of traditional methods, leading to suboptimal decisions.
In recent years, certain methods have been developed that account for interference in settings where only the quality of the final output is of interest (e.g., Ugander et al., 2013; Athey et al., 2018). The cumulative performance of the experimentation, which is particularly relevant in most online settings, has received relatively less attention, in part because it is more difficult to analyze. We aim to advance this line of research by studying the MAB problem in the presence of arbitrary network interference. Specifically, we consider $N$ units representing entities such as users on an online platform or a medical trial. There are $k$ treatment arms (or arms) and a time horizon of $T$ rounds. At each round, the learner assigns one arm to each unit and observes the resulting reward, with the goal of maximizing the total accumulated reward over all rounds. Unlike the traditional MAB framework, where rewards are assumed to be independent, interference introduces a dependency structure, with the reward of each unit determined not only by its own arm but also by the arms assigned to others. This dependency can significantly increase the complexity of exploration, as the action space may include up to $k^{N}$ different actions, growing exponentially in the number of units.
Classical MAB algorithms, such as the Upper Confidence Bound (UCB) method (Auer, 2002), yield a regret scaling of $\tilde{\mathcal{O}}(\sqrt{\frac{T}{N}k^{N}})$ [^1] Assuming each unit’s reward is 1-sub-Gaussian, the average reward over $N$ units follows a $(1/\sqrt{N})$-sub-Gaussian distribution., which becomes prohibitive as $N$ grows. Furthermore, without imposing any assumptions on the interference structure, the regret is guaranteed to scale as $\Omega(\sqrt{\frac{T}{N}k^{N}})$, as shown by lower bounds in the MAB literature (Lattimore & Szepesvári, 2020).
To address this challenge, we consider a setting where the reward of any given unit is influenced by its own arm and the arms assigned to its neighboring units, as defined by a graph that represents the interferences among the units. This graph is referred to as the interference graph.
Contributions Our main contributions are as follows.
  • We introduce the Partitioned Upper Confidence Bound with Interference (PUCB-I) algorithm for the MAB problem in the presence of network interference. Using the local structure of the interference graph, we establish a graph-dependent upper bound on cumulative regret. The algorithm outperforms state-of-the-art methods in all of the settings.
  • We derive the first lower bounds on the regret for bandits with arbitrary network interference. We provide two distinct lower bounds that quantify the worst-case regret of any algorithm based on the topological properties of the underlying interference graph. Specifically, these properties pertain to i) the symmetries in neighborhoods of the nodes (representing the units) and ii) the structure of doubly-independent sets of the interference graph. We demonstrate that the upper bound of our algorithm matches these lower bounds up to logarithmic factors for both dense and sparse graphs, thereby establishing its near-optimality. Furthermore, in general, the gap between the upper and lower bounds is at most $\sqrt{N}$.
  • Through numerical simulations, we show that the proposed algorithm outperforms the baseline approaches.
Paper Outline Following a review of related work, we provide an overview of the necessary background and key definitions in Section 2. In Section 3, we present our algorithm for minimizing cumulative regret and derive an upper bound on its regret. In Section 4, we establish theoretical lower bounds on regret. Finally, in Section 5, we demonstrate the efficiency of our algorithm through numerical simulations and showcase its advantage over the state of the art. Due to space limitations, some proofs are provided in Appendix A.
1.1 Related Work
Causal Inference and Interference.
The study of causal inference in the presence of interference, where the outcome of one unit is influenced by the treatments assigned to others, has been the focus of recent work in causal inference. Cross-unit interference violates the standard SUTVA (Rubin, 1980), a commonly made assumption in many traditional methods of experimental design and analysis. This problem has been studied in various fields such as statistics (Hudgens & Halloran, 2008; Eckles et al., 2017; Basse & Feller, 2018; Li & Wager, 2022; Leung, 2023), computer science (Ugander et al., 2013; Yuan et al., 2021; Ugander & Yin, 2023) and medical research (Tchetgen & VanderWeele, 2012). To address the challenge of interfering units, researchers have proposed tailor-made methodologies to estimate causal effects in structured interference models of their specific settings, such as intragroup interference (Rosenbaum, 2007; Hudgens & Halloran, 2008), network neighborhoods (Ugander et al., 2013; Bhattacharya et al., 2020; Yu et al., 2022; Gao & Ding, 2023), bipartite graphs (Pouget-Abadie et al., 2019; Bajari et al., 2021, 2023) and general graph models via exposure mappings (Aronow, 2012; Aronow & Samii, 2017).
Despite this extensive literature, all aforementioned works have focused merely on analyzing strategies that maximize the final reward. The harder problem of understanding cumulative performance over time remains relatively unexplored. The MAB framework is naturally well-suited to address this gap by balancing exploration and exploitation. Since the reward function in the setting of our interest is defined on $[k]^{N}$, the problem shares similarities with combinatorial bandits (Cesa-Bianchi & Lugosi, 2012), where the action space is defined as a subset of the binary hypercube. However, most prior work in the latter area has focused on linear reward functions. Recent efforts have extended combinatorial bandits to accommodate non-linear reward functions. However, these approaches often assume interference-free settings (Kveton et al., 2015; Agrawal et al., 2017) or focus on adversarial environments with restricted reward structures, for polynomial link functions (Han et al., 2021). Similarly, multiple-play bandits (Anantharam et al., 1987; Chen et al., 2013; Lagrée et al., 2016; Jia et al., 2023) consider settings where the learner selects multiple arms simultaneously and observes feedback for each arm. However, they also assume that the rewards are independent across arms, which is not the case in the presence of interference.
Recently, Jia et al. (2024) studied a batched adversarial bandit framework where $N$ units lie on the $\sqrt{N}\times\sqrt{N}$ unit grid. By limiting the action space to those with an identical arm for all units, they achieved a regret bound that does not depend exponentially on $N$. However, such an approach is quite limiting in practice, as the optimal action may be one that assigns heterogeneous arms to units.
A common approach to addressing the curse of dimensionality is to impose sparsity constraints, where only a fraction of actions yield non-zero rewards. For example, Hao et al. (2020) explored sparse linear bandits using an explore-then-commit algorithm, where actions are uniformly explored before applying Lasso regression to estimate the sparse structure of the reward parameters. Building on these approaches, Agarwal et al. (2024) studied a stochastic MAB problem with a similar motivation. Their work incorporated interference by defining the reward function over the hypercube $[k]^{N}$ and assuming that the reward of each unit is influenced by the arms assigned to its immediate neighbors, as well as its own arm. To address the curse of dimensionality, Agarwal et al. (2024) proposed a sparse network interference model, where each unit’s reward depends on the arms of at most $\Delta$ neighboring units. Using tools from discrete Fourier analysis, they developed a sparse linear representation of the reward function, enabling efficient computation. Their regression-based algorithm achieves a regret bound of $\tilde{\mathcal{O}}\left(T^{2/3}\right)$ which is worse than the classical MAB settings, highlighting the challenge posed by network interference. Agarwal et al. (2024) further proposed an alternative sequential action elimination algorithm that achieves regret proportional to $\sqrt{T}$, albeit growing with $N$, the number of units. In this paper, we propose the Partitioned UCB with Interference algorithm, which explicitly accounts for network interference by partitioning units based on their neighborhood structures in the interference graph. We establish an upper bound on regret that improves over previous work by a factor between $\sqrt{N}$ and $N$, depending on the interference graph. In particular, we avoid regrets growing with $N$, in contrast to previous work. Furthermore, we derive matching lower bounds (up to a logarithmic factor) and prove that our algorithm is near-optimal for both sparse and dense graphs.
2 Problem Setup
We consider a stochastic multi-armed bandit setting with $N$ units and $k$ available treatment arms (or simply, arms). At each round $t\in[T]$, an arm $A_{t}\in[k]^{N}$ is selected where $A_{ti}$ denotes the treatment assigned to unit $i\in[N]$. The reward function $Y_{ti}:[k]^{N}\to\mathbb{R}$ denotes the reward of the unit $i$, which depends on both its own treatment assignment and the treatment assignments of its neighbors due to interference.
To formally model the interference, we introduce an interference graph $\mathcal{G}\coloneqq([N],\mathcal{E})$ where the nodes $[N]$ represent the units, and an edge $(i,j)\in\mathcal{E}$ indicates that treatments assigned to $i$ and $j$ affect each other’s rewards.
The set of neighbors of a unit $i$ is denoted by $N(i)$, which includes all units $j$ connected to $i$ as well as $i$ itself. The size of this neighborhood is $|N(i)|=d_{i}+1$. At each round $t$, once treatment $A_{t}$ is assigned to the units, the learner observes the reward $Y_{ti}(A_{t})$  [^2] We assume that for each treatment $A$ and each unit $i$, the reward distribution $Y_{i}(A)$ is $1$-sub-Gaussian, a standard assumption in the bandit literature (Lattimore & Szepesvári, 2020; Bubeck et al., 2012). for each unit $i\in[N]$.
The assignment of treatments in each round is determined based on all previous treatment assignments and rewards. Formally, a policy $\pi\coloneqq(\pi_{1},\dots,\pi_{T})$ represents a sequence of adaptive mappings such that:
$\pi_{t}:\big{(}[k]^{N}\times\mathbb{R}^{N}\big{)}^{t-1}\to\mathcal{P}([k]^{N}),$
where $\mathcal{P}([k]^{N})$ denotes the space of probability distributions over the treatment assignments $[k]^{N}$, and $\big{(}[k]^{N}\times\mathbb{R}^{N}\big{)}^{t-1}$ represents the accumulated history of treatment assignments and rewards up to time $t-1$.
At each round $t$, the treatment assignment $A_{t}$ is sampled from the policy as:
$A_{t}\sim\pi_{t}\left(A_{1},Y_{1},\ldots,A_{t-1},Y_{t-1}\right).$
To evaluate the performance of a treatment assignment policy over $T$ rounds, we define cumulative regret as the gap between the optimal reward in hindsight and the expected reward achieved by the policy.
The expected reward for a treatment $A$ for unit $i$ is defined using a reward function $\mu_{i}$, given by:
$\mu_{i}(A)\coloneqq\mathbbm{E}[Y_{i}|A_{N(i)}].$
For any unit $i\in[N]$ and any treatment $A\in k^{N}$, we assume $\mu_{i}(A)\in[0,1]$.
Definition 2.1 (Regret).
The regret of policy $\pi$ that interacts with an instance $\mathcal{V}$ is defined as:
$Reg_{T}(\pi,\mathcal{V})\coloneqq$ $\frac{1}{N}\mathbb{E}_{A_{t}\sim\pi}\Big{[}\max_{A\in[k]^{N}}T\sum_{i\in[N]}\mu_{i}(A)-\sum_{\begin{subarray}{c}i\in[N]\\ t\in[T]\end{subarray}}\mu_{i}(A_{t})\Big{]},$
where the first term represents the cumulative reward achievable under an optimal treatment assignment, maximizing the expected reward and the second term corresponds to the cumulative reward achieved by the policy $\pi$ through its treatment assignments $A_{t}$ at each round. Here, $\mathcal{V}$ denotes the instance, including the interference graph and the reward distributions, which determine how the rewards depend on the treatment assignments. For simplicity, we may drop $\pi$ or $\mathcal{V}$ from $Reg_{T}(\pi,\mathcal{V})$ when they are clear from the context.
Remark 2.2.
Note that if the interference graph $\mathcal{G}$ is not connected, each of its connected components can be treated as an independent instance. The problem can then be solved separately for each component and the total regret of any algorithm can be expressed as the sum of the regrets over the individual connected components. The same principle applies to proving the lower bounds. Therefore, for the remainder of this paper, we assume that the interference graph $\mathcal{G}$ is connected.
3 Upper bound
In this section, we propose the Partitioned UCB with Interference (PUCB-I) algorithm to address the MAB problem with network interference, where the reward of each unit depends on its treatment assignment and the treatment assignments of its neighbors. The algorithm relies on knowledge of the interference graph structure to balance exploration and exploitation, ensuring efficient learning.
First, we define a relation on the pairs of units which leads to a partitioning of units that PUCB-I uses to choose its next treatment assignment.
Define a binary relation on a pair of units $i,i^{\prime}\in[N]$, such that $i\sim i^{\prime}$ if they are connected and have identical neighborhoods in the interference graph $\mathcal{G}$, that is,
$i\sim i^{\prime}\Longleftrightarrow N(i)=N(i^{\prime}).$ (1)
It is straightforward to verify that this is an equivalence relation and partitions $[N]$ into $M$ equivalence classes $\{P_{1},P_{2},\dots,P_{M}\}$. For each partition, denote its size by $m_{j}=|P_{j}|$. In this partitioning, all nodes in a partition $P_{j}$ have identical degrees since they share the same neighborhood. We denote this common degree for the nodes in partition $P_{j}$ as $D_{j}$. Figure 1 depicts one such partitioning. It can be easily verified that in this example the nodes are partitioned into $P_{1}=\{A,B\}$, $P_{2}=\{C\}$, $P_{3}=\{D,E\}$, $P_{4}=\{F,G,H\}$. For a graph $\mathcal{G}$, we denote by $M(\mathcal{G})$ the number of partitions of $\mathcal{G}$ with respect to the relation in Equation (1). When the graph is clear from the context, we simply use $M$ instead of $M(\mathcal{G})$.
ABCDEFGH
Figure 1: Partitions of a graph with $8$ nodes according to the equivalence relation defined in Eq. (1). The nodes in each partition are shown in the same color.
Now, we introduce the key components of our algorithm.
Initialization.
The algorithm begins with an initial exploration phase, ensuring that every possible treatment $A_{N(i)}$ is assigned at least once for each unit $i$. Since units in the same partition $P_{j}$ share identical neighborhood structures, the algorithm does not need to perform an exploration for each unit separately. Instead, covering all possible treatments for partition $P_{j}$ requires $k^{D_{j}+1}$ rounds. Therefore, the total number of exploration rounds needed across all partitions is $\sum_{j\in[M]}k^{D_{j}+1}$.
Empirical Mean Reward.
For each unit $i\in[N]$ and treatment $A_{N(i)}\in[k]^{d_{i}+1}$, the empirical mean reward is estimated as:
$\hat{\mu}_{ti}(A)=\frac{\sum_{t^{\prime}=1}^{t}Y_{t^{\prime}i}(A_{N(i)})}{n_{ti}(A)},$
where $n_{ti}(A)$ is the number of times that $A_{N(i)}$ has been assigned to unit $i$ up to time $t$, or
$n_{ti}(A)=\sum_{t^{\prime}=1}^{t}\mathbbm{1}\{A_{t^{\prime}N(i)}=A_{N(i)}\}.$
By definition of our partitioning, for every treatment $A\in[k]^{N}$ and $i,i^{\prime}\in P_{j}$, we have $n_{ti}(A)=n_{ti^{\prime}}(A)$. Therefore, we define $n_{tP_{j}}(A)\coloneqq n_{ti}(A)$ which holds for every unit $i\in P_{j}$.
UCB. For each partition $P_{j}$ at round $t$, compute the UCB for every treatment $A$ as:
$UCB_{tP_{j}}(A)\coloneqq\sum_{i\in P_{j}}\hat{\mu}_{ti}(A)+\sqrt{2\log(2/\delta)\frac{m_{j}}{n_{tP_{j}}(A)}}.$ (2)
Treatment Assignment. The treatment assignment for the next round is determined by maximizing the sum of UCBs across all partitions:
$A_{t+1}=\operatorname*{arg\,max}_{A\in[k]^{N}}\sum_{j\in[M]}UCB_{tj}(A).$ (3)
Reward Observation and Update. After selecting $A_{t+1}$, the algorithm observes the rewards $Y_{t+1i}(A_{t+1})$ for all $i\in[N]$, and updates the empirical mean reward estimates $\hat{\mu}_{t+1i}(A)$ and the count $n_{t+1P_{j}}(A)$ for the explored treatments.
Algorithm 1 Partitioned UCB with Interference (PUCB-I)
1:  Input: Number of rounds $T$, interference graph $\mathcal{G}$, number of treatment arms $k$, confidence parameter $\delta$.
2:  Initialization: Collect one sample from each possible treatments $A_{N(i)}$ for each unit $i\in[N]$.
3:  for each round $t=\sum_{j\in[M]}k^{D_{j}+1},\dots,T-1$ do
4:     for each partition $P_{j}\in\{P_{1},\dots,P_{M}\}$ do
5:        Compute $\text{UCB}_{tP_{j}}(A)$ for all $A\in[k]^{N}$ using Eq. (2).
6:     end for
7:     Select $A_{t+1}$ that maximizes $\sum_{j\in[M]}\text{UCB}_{tP_{j}}(A)$ using Eq. (3).
8:     Observe rewards $Y_{t+1i}(A_{t+1})$ for all $i\in[N]$.
9:     Update $\hat{\mu}_{t+1i}(A)$ and $n_{t+1P_{j}}(A)$ for the explored treatments.
10:  end for
The following theorem establishes a graph-dependent upper bound on the expected cumulative regret of Algorithm 1:
{restatable}
theoremthupper[Graph-Partitioned Regret Upper Bound] The expected cumulative regret of Algorithm 1 with $\delta=(T^{2}N\sum_{j\in[M]}k^{D_{j}+1})^{-1}$ interacting with any bandit instance with $1$-sub-Gaussian reward distributions and interference graph $\mathcal{G}$ with partitions $P_{1},P_{2},\ldots,P_{M}$ over $T$ rounds satisfies:
$Reg_{T}\in\mathcal{O}\left(\sqrt{\frac{T}{N^{2}}\log(TN)}\sum_{j\in[M]}\sqrt{m_{j}k^{D_{j}+1}}\right).$
Proof Sketch. Here, we outline the main steps of the proof. The complete proof is provided in Appendix A.
First, we define a good event $G$ as
$G\coloneqq\mathbbm{1}$$\Bigg{\{}\forall t\in[T],j\in[M],A\in[k]^{N}:$
$\left|\sum_{i\in P_{j}}\hat{\mu}_{ti}(A)-\mu_{ti}(A)\right|\leq\sqrt{2\log(2/\delta)\frac{m_{j}}{n_{tP_{j}}(A)}}\Bigg{\}}.$
When $G$ holds, for each partition and treatment within a partition, the true sum of rewards remains within the confidence interval around the empirical sum for all rounds $t\in[T]$. By applying Hoeffding’s inequality and the union bound, we can bound the probability of $G^{c}$ as follows:
$P(G^{c})\leq T\sum_{j\in[M]}k^{D_{j}+1}\delta.$
On the other hand, if $G$ holds, we can bound the regret of PUCB-I as:
${Reg}_{T}\leq\frac{1}{N}\left(2\sqrt{2\log(2/\delta)}\sum_{t\in[T]}\sum_{j\in[M]}\sqrt{\frac{m_{j}}{n_{tP_{j}}(A_{t})}}\right).$
The next step is to bound the sum $\sum_{t\in[T]}\sum_{j\in[M]}\sqrt{\frac{m_{j}}{n_{tP_{j}}(A_{t})}}$. Using the fact that for units in $P_{j}$, the number of different treatments is $k^{D_{j}+1}$, and applying the inequality $\forall n\in\mathbb{N}:\sum_{i\in[n]}\frac{1}{\sqrt{i}}\leq 2\sqrt{n}$, we obtain:
$\sum_{t\in[T]}\sum_{j\in[M]}\sqrt{\frac{m_{j}}{n_{tP_{j}}(A_{t})}}\leq 2\sum_{j\in[M]}\sqrt{Tm_{j}k^{D_{j}+1}}.$
Then, conditioning the value of regret on the event $G$ and using the assumption that for all $i\in[N],A\in k^{N}$ it holds $\mu_{i}(A)\in[0,1]$, we obtain a bound on the regret as:
${Reg}_{T}\leq\frac{1}{N}$ $\Bigg{(}4\sqrt{2\log(2/\delta)T}\sum_{j\in[M]}\sqrt{m_{j}k^{D_{j+1}}}$
$+TN\bigg{(}T\sum_{j\in[M]}k^{D_{j}+1}\delta\bigg{)}\Bigg{)}.$
Finally, setting $\delta=(T^{2}N\sum_{j\in[M]}k^{D_{j}+1})^{-1}$ and using $T\geq k^{\Delta+1}$, the proof is complete. ∎
The regret scales as $\mathcal{O}(\sqrt{T})$, as a result of the interplay between exploration and exploitation over $T$ rounds. The dependency on $k^{D_{j}+1}$ reflects the complexity introduced by interference through the size of the neighborhood in each partition $P_{j}$. A perhaps more interpretable form of the bound can be obtained in terms of the maximum degree, $\Delta:=\max_{i\in[N]}d_{i}$, as follows:
Corollary 3.1.
Let $\Delta$ denote the maximum degree of $\mathcal{G}$. Then, the expected cumulative regret of Algorithm 1 up to time $T$ is bounded as:
$Reg_{T}\in\mathcal{O}\left(\sqrt{\frac{TM}{N}k^{\Delta+1}\log(TN)}\right).$
The corollary follows from applying the Cauchy-Schwarz inequality:
$\sum_{j\in[M]}\sqrt{m_{j}k^{\Delta+1}}\leq\sqrt{M\sum_{j\in[M]}m_{j}k^{\Delta+1}}=\sqrt{MNk^{\Delta+1}}.$
Corollary 3.1 allows us to compare the regret of our algorithm versus the state-of-the-art. In Agarwal et al. (2024), the authors established a regret upper bound of $\tilde{\mathcal{O}}(\sqrt{TNk^{\Delta+1}})$, which is $N/\sqrt{M}$ worse than the regret bound of PUCB-I. At one end, when $M=N$ (i.e., each unit forms its own partition), the gap is $\sqrt{N}$, while at the other end, when $M=1$, it can be as large as $N$.
If the proposed partitioning were ignored and the upper confidence bounds were computed for each unit—treating each unit as its own partition—the resulting regret upper bound would be
$\mathcal{O}\left(\sqrt{\frac{T}{N^{2}}\log(TN)}\sum_{i\in[N]}\sqrt{k^{d_{i}+1}}\right).$
This bound is always at least as large as the bound in Theorem 1. This follows from the inequality
$\sum_{i\in[N]}\sqrt{k^{d_{i}+1}}=\sum_{j\in[M]}m_{j}\sqrt{k^{D_{j}+1}}\geq\sum_{j\in[M]}\sqrt{m_{j}k^{D_{j}+1}}.$
This indicates that partitioning reduces the total regret induced by the units in partition $P_{j}$ by a factor of $\sqrt{m_{j}}$. For example, in the case of a complete graph, this leads to a $\sqrt{N}$ improvement in regret.
4 Lower Bounds
In this section, we establish the first lower bounds on the expected regret for MABs with arbitrary network interference.
We derive two distinct bounds that quantify the worst-case regret of any algorithm based on the topological properties of the underlying interference graph. Specifically, these properties pertain to i) the symmetries in neighborhoods of the nodes and ii) the structure of doubly-independent sets of the interference graph.
Both lower bounds have a gap with the proposed upper bound for PUCB-I in Theorem 1. Subsequently, we identify classes of graphs where these gaps are constant, showing that our algorithm is nearly optimal (up to logarithmic factors). The first lower bound demonstrates that PUCB-I is near-optimal for classes of dense graphs, while the second proves that it is near-optimal for sparse graphs. Sparse graphs are particularly significant in multi-armed bandit problems with interference, as many practical applications involve interference graphs that are sparse due to limited local interactions (Agarwal et al., 2024; Yang et al., 2016).
The following theorem establishes our first lower bound on the expected regret.
{restatable}
theoremtheoremlowerbound [Graph-Partitioned Regret Lower Bound] Let $\mathcal{G}$ be a connected graph with $N$ nodes and partitions $\left\{P_{1},P_{2},\ldots,P_{M}\right\}$ with $|P_{j}|=m_{j}$. If $k>\frac{2^{\frac{1}{\Delta+1}}}{2^{\frac{1}{\Delta+1}}-1}$ and $T\geq\frac{4(k-1)^{\Delta+1}}{M}$, then for any policy $\pi$, there exists a bandit instance with interference graph $\mathcal{G}$ whose reward is distributed as $1$-Gaussian with means in $[0,1]$ such that
$Reg_{T}(\pi)\in\Omega\left(\sqrt{\frac{T}{N^{2}M}}\sum_{j\in[M]}\sqrt{m_{j}k^{D_{j}+1}}\right).$
This lower bound has a gap of $\sqrt{M}$ compared to the upper bound of PUCB-I. This indicates that the algorithm achieves better performance for graphs with a smaller number of partitions.
As discussed earlier, interference graphs in real-world scenarios are often sparse and have relatively few edges. We now derive a second lower bound on the expected regret that nearly matches the upper bound of our proposed algorithm for graphs with a bounded maximum degree. To begin, we introduce two essential definitions.
Definition 4.1 (Doubly-Independent Set).
Let $\mathcal{G}$ be a connected graph with $N$ nodes. A set of nodes $S$ is called a doubly-independent set if no two nodes in $S$ are adjacent, nor do they share a common neighbor outside $S$. The family of all doubly-independent sets of $\mathcal{G}$ is denoted by $DI(\mathcal{G})$.
Definition 4.2 (Square Chromatic Number).
The square chromatic number of a graph $\mathcal{G}$ is the minimum number of colors required to color its nodes such that all nodes of the same color form a doubly-independent set. The square chromatic number of $\mathcal{G}$ is denoted by $\chi(\mathcal{G}^{2})$.
Theorem 4.3 (Doubly-Independent Set Regret Lower Bound).
Let $\mathcal{G}$ be a connected graph with $N$ nodes. If $T\geq k^{\Delta+1}-1$, for any policy $\pi$, there exists a bandit instance with interference graph $\mathcal{G}$, $1$-Gaussian reward distributions and means in $[0,1]$ such that
$Reg_{T}(\pi)\in\Omega\left(\max_{S\in DI(\mathcal{G})}\sqrt{\frac{T}{N^{2}}}\sum_{i\in S}\sqrt{k^{d_{i}+1}}\right).$ (4)
Proof.
To prove this theorem, we first provide two important lemmas.
{restatable}
lemmaedgeremoval For a graph $\mathcal{G}$, let $\mathcal{G}^{\prime}$ be the graph obtained after removing an arbitrary edge from $\mathcal{G}$. Then, for any policy $\pi$, the worst-case regret of interacting with bandit instances on $\mathcal{G}$ is at least that of interacting with bandit instances on $\mathcal{G}^{\prime}$. That is,
$\sup_{\mathcal{V}\sim\mathcal{G}}Reg_{T}(\pi,\mathcal{V})\geq\sup_{\mathcal{V}^{\prime}\sim\mathcal{G}^{\prime}}Reg_{T}(\pi,\mathcal{V}^{\prime}),$
where $V\sim\mathcal{G}$ denotes a bandit instance whose interference graph is $\mathcal{G}$.
{restatable}
lemmaonenodelowerbound Let $\mathcal{G}$ be a graph with $N$ nodes where $d_{i}$ is the degree of node $i$. If $T\geq k^{d_{i}+1}-1$, for any policy $\pi$ and each $i\in[N]$, there exists a bandit instance with interference graph $\mathcal{G}$, $1$-Gaussian reward distributions and means in $[0,1]$ such that:
$Reg_{T}(\pi)\in\Omega\left(\sqrt{\frac{Tk^{d_{i}+1}}{N^{2}}}\right).$
We prove the theorem using the aforementioned lemmas as follows. For any set $S=\{s_{1},s_{2},\ldots,s_{m}\}\in DI(\mathcal{G})$, let $\mathcal{G}_{S}$ be the graph obtained from $\mathcal{G}$ by removing all edges between nodes that are outside of $S$ (i.e., edges where neither endpoint belongs to $S$). By Lemma 4, the worst-case regret for instances with interference graph $\mathcal{G}$ is at least as large as the worst-case regret for instances with $\mathcal{G}_{S}$. The graph $\mathcal{G}_{S}$ consists of edges where at least one endpoint lies in $S$. Since $S\in DI(\mathcal{G})$, the resulting graph $\mathcal{G}_{S}$ consists of $m$ disjoint connected components $\mathcal{G}_{s_{1}},\mathcal{G}_{s_{2}}\ldots,\mathcal{G}_{s_{m}}$. Each component $\mathcal{G}_{s_{i}}$ is a star graph with $d_{s_{i}}+1$ nodes.
By Lemma 4, for any policy $\pi$ interacting with $\mathcal{G}_{s_{i}}$, there exists a bandit instance where:
$Reg_{T}(\pi)\in\Omega\left(\sqrt{\frac{Tk^{d_{s_{i}}+1}}{N^{2}}}\right).$
Since the components $\mathcal{G}_{s_{i}}$s are disjoint, the total regret for the entire graph $\mathcal{G}_{S}$ is the sum of the regrets for each component. Therefore, for any policy $\pi$ and any $S\in DI(\mathcal{G})$, there exists an instance such that:
$Reg_{T}(\pi)\in\Omega\left(\sqrt{\frac{T}{N^{2}}}\sum_{i\in S}\sqrt{k^{d_{i}+1}}\right).$
This concludes the proof.
The following corollary establishes the connection between doubly-independent sets and the square chromatic number of a graph, enabling a comparison between the two lower bounds.
Corollary 4.4.
A connected graph $\mathcal{G}$ can be colored using $\chi(\mathcal{G}^{2})$ distinct colors where each color class forms a doubly-independent set. Therefore, the maximum in (4) is at least as large as any such set, or
$Reg_{T}(\pi)\in\Omega\left(\frac{1}{\chi(\mathcal{G}^{2})}\sqrt{\frac{T}{N^{2}}}\sum_{i\in[N]}\sqrt{k^{d_{i}+1}}\right).$ (5)
The first lower bound has a $\sqrt{M(\mathcal{G})}$ gap with the upper bound. This gap for the second lower bound is $\chi(\mathcal{G}^{2})$. Consequently, when $\chi(\mathcal{G}^{2})>\sqrt{M(\mathcal{G})}$, the first bound is tighter, and vice versa. Overall, the gap between the upper and lower bounds is at most $\min\left(\sqrt{M(\mathcal{G})},\chi(\mathcal{G}^{2})\right)$ which is always bounded by $\sqrt{N}$. Therefore, for graphs with constant $\sqrt{M(\mathcal{G})}$ or $\chi(\mathcal{G}^{2})$, PUCB-I achieves near-optimal regret.
4.1 Graphs with Tight Bounds
This section explores the classes of graphs where $\min\left(\sqrt{M(\mathcal{G})},\chi(\mathcal{G}^{2})\right)$ is constant.
To identify the graphs for which the first lower bound demonstrates the near-optimality of our algorithm, that is, $\sqrt{M(\mathcal{G})}$ is constant, we define a class of dense graphs, which we call Clique-Sparse Graphs.
Definition 4.5 (Clique-Sparse Graph).
A graph $\mathcal{G}$ with $N$ nodes is called $(R,r)$-Clique-Sparse if it is possible to partition the nodes into $R$ clusters $C_{1},C_{2},\ldots,C_{R}$ satisfying the following conditions:
  • $\forall i\in[R]$, cluster $C_{i}$ forms a complete graph.
  • The number of edges between the nodes in any pair of clusters is at most $r$.
ABCDEFGHIJK
Figure 2: A $(3,2)$-clique-sparse graph with 11 Nodes. The nodes in each cluster are shown in the same color.
As an example, note that the graph in Figure 2 is $(3,2)$-clique-sparse, and the complete graph is $(1,0)$-clique-sparse.
The following lemma shows that for $(R,r)$-clique-sparse graphs $\mathcal{G}$, where both $R$ and $r$ are constant, the number of partitions $M(\mathcal{G})$ is also constant. This result establishes that PUCB-I is near-optimal for this class of graphs.
{restatable}
lemmacliquesparse For a $(R,r)$-clique-sparse graph $\mathcal{G}$, the number of partitions $M(\mathcal{G})$ induced by the equivalence relation in Equation (1) satisfies:
$M(\mathcal{G})\leq R+rR(R-1).$
The following lemma shows that for sparse graphs with a constant maximum degree $\Delta$, the square chromatic number is also a constant which implies that PUCB-I is near optimal for such graphs.
{restatable}
lemmaconstantchrom For any graph $\mathcal{G}$ with maximum degree $\Delta$, the square chromatic number $\chi(\mathcal{G}^{2})$ is bounded as
$\chi(\mathcal{G}^{2})\leq\Delta^{2}+1.$
5 Experiments
In this section, we perform simulations to empirically evaluate the performance of our algorithm and validate its theoretical guarantees. We compare the results of Algorithm 1 (PUCB-I) with three other baselines:
  • Classical UCB. This algorithm ignores the interference graph, treats each treatment in $k^{[N]}$ as an independent arm, and performs the UCB algorithm.
  • Network Explore-Then-Commit (ETC). This algorithm, introduced by Agarwal et al. (2024), operates in two phases. First, it assigns treatments uniformly at random for an initial phase. It then estimates the reward parameters using least squares regression, incorporating the known interference graph. In the second phase, it selects and plays the arm with the highest estimated reward for the remaining rounds.
  • Sequential Action Elimination (SAE). Also introduced by Agarwal et al. (2024), this elimination-based algorithm follows an epoch-based structure. It begins with all $k^{[N]}$ arms considered active. In each epoch, it pulls all active arms an equal number of times and progressively eliminates those with poor observed performance.
The complete implementation details are provided in the supplementary material. In all experiments, we set $k=2$, the total number of rounds $T_{max}=10\cdot 2^{N}$, to ensure sufficient exploration of all $2^{N}$ arms as required by the classical UCB algorithm. Additionally, the rewards for each unit were randomly drawn from a $1$-Guassian distribution. In two sets of experiments, we analyzed the cumulative regret as a function of $T$ and $N$. The results were averaged over $50$ independent iterations.
Scaling of Regret with $T$.
Figure 3 presents the cumulative regret of different algorithms over time horizon $T$ for the number of units $N=10$ and the maximum degree $\Delta=4$. As expected, classical UCB requires extensive exploration to converge, while our algorithm achieves significantly lower regret by partitioning units based on their neighborhoods, reducing unnecessary exploration. ETC also improves upon UCB but requires longer exploration phases.
Figure 3: A comparison of the average regret of various algorithms over time ($T$).
Scaling of Regret with $N$.
To evaluate the impact of the number of units $N$, we analyzed the cumulative regret of the algorithms across instances by varying $N$ in $\{4,6,8,10,12\}$. For all the instances, we set the maximum degree $\Delta=\frac{N}{2}-1$. As shown in Figure 4, when $N$ increases, the regret of UCB grows exponentially, confirming that it was not able to incorporate the graph structure. In contrast, the regret of Algorithm 1 (PUCB-I) grows slower as $N$ increases compared to the baselines because it benefits from partitioning the units with similar neighborhoods.
Figure 4: A comparison of the average regret of various algorithms for different numbers of units ($N$).
6 Conclusion
We studied the MAB problem with network interference, where the reward of each unit depends on both its assigned treatment and the treatments of its neighbors. To address the challenges posed by interference, we proposed the PUCB-I algorithm, which partitions units based on their neighborhood structure to minimize cumulative regret. We established a graph-dependent regret upper bound and derived the first regret lower bounds for bandits with arbitrary network interference, showing that our proposed algorithm achieves near-optimal performance (up to logarithmic factors) for both sparse and dense graphs. Our work highlights the importance of accounting for the graph structure in sequential decision-making under interference. Empirical results further validate our theoretical results, showing that PUCB-I outperforms existing baselines.
Acknowledgments
This research was in part supported by the Swiss National Science Foundation under NCCR Automation, grant agreement 51NF40_180545.
Appendix A Omitted Proofs
[…]
Proof.
By the definition of $A_{t+1}$, we have:
$\sum_{i\in[N]}\hat{\mu}_{ti}(A_{t})+\sum_{j\in[M]}\sqrt{2\log(2/\delta)\frac{m_{j}}{n_{tP_{j}}(A_{t})}}\geq\sum_{i\in[N]}\hat{\mu}_{ti}(A^{*})+\sum_{j\in[M]}\sqrt{2\log(2/\delta)\frac{m_{j}}{n_{tP_{j}}(A^{*})}},$ (6)
where $A^{*}\in\operatorname*{arg\,max}_{A\in[k]^{N}}\sum_{i\in[N]}\mu_{i}(A)$ denotes the optimal treatment. We define the good event $G$, ensuring that the empirical mean reward is close to the true mean for all times $t$, partitions $j$, and treatments $A$:
$G\coloneqq\mathbbm{1}\left\{\forall t\in[T],j\in[M],A\in[k]^{N}:\left|\sum_{i\in P_{j}}\hat{\mu}_{ti}(A)-\mu_{ti}(A)\right|\leq\sqrt{2\log(2/\delta)\frac{m_{j}}{n_{tP_{j}}(A)}}\right\}.$
Applying the Hoeffding’s inequality, we obtain:
$P\left(\left|\sum_{i\in P_{j}}\hat{\mu}_{ti}(A)-\mu_{ti}(A)\right|\geq\sqrt{2\log(2/\delta)\frac{m_{j}}{n_{tP_{j}}(A)}}\right)\leq\delta.$
By a union bound over all $t$, $j$ and $A$:
$P(G^{c})$ $\leq\sum_{t\in[T]}\sum_{j\in[M]}\sum_{A\in[k]^{N}}P\left(\left|\sum_{i\in P_{j}}\hat{\mu}_{ti}(A)-\mu_{ti}(A)\right|\geq\sqrt{2\log(2/\delta){\frac{m_{j}}{n_{tP_{j}}(A)}}}\right)$
$\leq T\sum_{j\in[M]}k^{D_{j}+1}\delta,$
where $D_{j}$ denotes the degree of each unit in the $j$-th partition.
By the law of total expectations, we can write the regret under event $G$ as:
$N.{Reg}_{T}$ $=\sum_{t\in[T]}\sum_{i\in[N]}\mu_{ti}(A^{*})-\mu_{ti}(A_{t})$
$\overset{(a)}{\leq}\sum_{t\in[T]}\sum_{i\in[N]}\hat{\mu}_{ti}(A^{*})+\sum_{j\in[M]}\sqrt{2\log(2/\delta)\frac{m_{j}}{n_{tP_{j}}(A^{*})}}-\left[\hat{\mu}_{ti}(A_{t})-\sum_{j\in[M]}\sqrt{2\log(2/\delta)\frac{m_{j}}{n_{tP_{j}}(A_{t})}}\right]$
$\overset{(b)}{\leq}2\sqrt{2\log(2/\delta)}\sum_{t\in[T]}\sum_{j\in[M]}\sqrt{\frac{m_{j}}{n_{tP_{j}}(A_{t})}}$
where $(a)$ and $(b)$ hold by definition of event $G$ and (6), respectively.
Next, we bound $\sum_{t\in[T]}\frac{1}{\sqrt{n_{tP_{j}}(A_{t})}}$ as
$\sum_{t\in[T]}\frac{1}{\sqrt{n_{tP_{j}}(A_{t})}}$ $=\sum_{A\in[k]^{D_{j}+1+1}}\sum_{t\in[T]}\mathbbm{1}\{\forall i\in P_{j}:A_{tN(i)}=A_{N(i)}\}\frac{1}{\sqrt{n_{tP_{j}}(A)}}$ (7)
$=\sum_{A\in[k]^{D_{j}+1}}\sum_{t\in[n_{TP_{j}}(A)]}\frac{1}{\sqrt{t}}$
$\overset{(a)}{\leq}2\sum_{A\in[k]^{D_{j}+1}}\sqrt{n_{TP_{j}}(A)}$
$\overset{(b)}{\leq}2\sqrt{k^{D_{j}+1}\sum_{A\in[k]^{D_{j}+1}}n_{TP_{j}}(A)}$
$\leq 2\sqrt{Tk^{D_{j}+1}}$
where $(a)$ holds since $\forall n\in\mathbb{N}:\sum_{i\in[n]}1/\sqrt{i}\leq 2\sqrt{n}$, and $(b)$ follows from Jensen inequality.
We prove the inequality $\sum_{i=1}^{n}\frac{1}{\sqrt{i}}\leq 2\sqrt{n}$ using integral approximation. Since the function $f(x)=\frac{1}{\sqrt{x}}$ is decreasing,
[…]
Evaluating the integral,
[…]
Thus,
$\sum_{i=1}^{n}\frac{1}{\sqrt{i}}\leq 1+(2\sqrt{n}-2)\leq 2\sqrt{n}.$
Therefore, the regret under event $G$ is:
$N.Reg_{T}\leq 4\sqrt{2\log(2/\delta)T}\sum_{j\in[M]}\sqrt{m_{j}k^{D_{j+1}}}$ (8)
Thus, by choosing $\delta=\frac{1}{T^{2}N\sum_{j\in[M]}k^{D_{j}+1}}$ we get the following by the law of total probability:
${Reg}_{T}$ $\leq 4/{N}\sqrt{2\log(2/\delta)T}\sum_{j\in[M]}\sqrt{m_{j}k^{D_{j+1}}}+T(T\sum_{j\in[M]}k^{D_{j}+1}\delta)$
$\in\mathcal{O}\left(\sqrt{\frac{T}{N^{2}}\log(T^{2}N\sum_{j\in[M]}k^{D_{j}+1})}\sum_{j\in[M]}\sqrt{m_{j}k^{D_{j+1}}}\right).$
Therefore, since $T\geq k^{\Delta+1}$, it satisfies:
${Reg}_{T}\in\mathcal{O}\left(\sqrt{\frac{T}{N^{2}}\log(TN)}\sum_{j\in[M]}\sqrt{m_{j}k^{D_{j+1}}}\right).$
[…]
Proof.
To prove this lower bound, we employ a change-of-measure argument, a well-known technique in the multi-armed bandit literature for deriving lower bounds. Consider an instance $\mathcal{V}$ with interference graph $\mathcal{G}$, Gaussian reward noises with unit variance, and the following reward means for each unit $i\in[N]$:
$\mu(A_{N(i)})=\begin{cases}\Delta_{i}&\text{if }A_{N(i)}=(1,1,\ldots,1),\\ 0&\text{otherwise,}\end{cases}$
where $\Delta_{i}$ is a positive real number to be determined later. In this case, the optimal treatment is to assign treatment $1$ to all units, yielding an expected reward of $\sum_{i\in[N]}\Delta_{i}$.
Now, fix a policy $\pi$ operating on instance $\mathcal{V}$. For each treatment $A\in[k]^{N}$, let $T_{\pi}(A)=\left(t_{\pi,1}(A),t_{\pi,2}(A),\ldots,t_{\pi,N}(A)\right)$, where $t_{\pi,i}(A)$ is the expected number of times policy $\pi$ applies a treatment assignment such that unit $i$ and its neighbors receive treatment $A_{N(i)}$ over $T$ rounds. For simplicity, we denote $T_{\pi}(A)$ and $t_{\pi,i}(A)$ as $T(A)$ and $t_{i}(A)$, respectively.
Let $S\subset[k]^{N}$ be the set of all treatments where no unit receives treatment $1$, therefore, $|S|=(k-1)^{N}$. The following lemma implies that there is always an under-explored treatment.
Lemma A.1.
For any policy $\pi$ and any set of values $\Delta_{i}$ for $i\in[N]$, there exists a treatment $A^{\prime}\in S$ such that:
$\sum_{i\in[N]}t_{i}(A^{\prime})\Delta_{i}^{2}\leq T\left(\sum_{i\in[N]}\frac{\Delta_{i}^{2}}{(k-1)^{d_{i}+1}}\right).$
Proof.
To prove this lemma, we employ a double-counting technique. Consider a matrix with $(k-1)^{N}$ rows and $N$ columns where row $i$ corresponds to a treatment $A_{i}\in S$ and column $i$ corresponds to unit $i$. On the element in $j$-th row and $i$-th column of the matrix, we write $t_{i}(A_{j})\Delta_{i}^{2}$ as follows:
$\mathbf{M}\coloneqq\begin{bmatrix}t_{1}(A_{1})\Delta_{1}^{2}&t_{2}(A_{1})\Delta_{2}^{2}&\cdots&t_{N}(A_{1})\Delta_{N}^{2}\\ t_{1}(A_{2})\Delta_{1}^{2}&t_{2}(A_{2})\Delta_{2}^{2}&\cdots&t_{N}(A_{2})\Delta_{N}^{2}\\ \vdots&\vdots&\ddots&\vdots\\ t_{1}(A_{k-1^{N}})\Delta_{1}^{2}&t_{2}(A_{k-1^{N}})\Delta_{2}^{2}&\cdots&t_{N}(A_{k-1^{N}})\Delta_{N}^{2}\end{bmatrix}.$
To calculate the sum of the numbers in the first column, we have
$\sum_{A_{i}\in S}t_{1}(A_{i})\Delta_{1}^{2}=\Delta_{1}^{2}(k-1)^{N-(d_{i}+1)}\sum_{A_{N(1)}\in S_{1}}t(A_{N(1)}),$
where $S_{1}=\{2,3,\dots,k\}^{d_{i}+1}$ is the set of all treatments that can be assigned to unit one and its neighbors without using treatment $1$ and $t(A_{N(1)})$ is the expected number of times that unit one and its neighbors are assigned $A_{N(1)}$ during $T$ rounds of interaction between $\pi$ and $\mathcal{V}$. The equation is true because of the symmetry in the problem which implies that each combination of the treatment of unit $1$ and its neighbors exists in $(k-1)^{N-(d_{i}+1)}$ number of members of $S$. On the other hand, we have
$\sum_{A_{N(1)}\in S_{1}}t(A_{N(1)})=T.$
Writing the same equation for all the columns implies that the sum of the numbers in the whole matrix is equal to
$T\sum_{i\in[N]}\Delta_{i}^{2}(k-1)^{N-(d_{i}+1)}.$
Dividing this number by the number of rows shows that there exists a row $j$ such that
$\sum_{i\in[N]}t_{i}(A_{j})\Delta_{i}^{2}\leq T\left(\sum_{i\in[N]}\frac{\Delta_{i}^{2}}{(k-1)^{d_{i}+1}}\right).$
Now, based on the under-explored treatment $A^{\prime}$, we design a confusing instance $\mathcal{V}^{\prime}$ which for each treatment $A$ has the mean expected rewards as follows:
$\mu^{\prime}(A_{N(i)})=\begin{cases}\Delta_{i}\quad&\text{ if }A_{N(i)}=(1,1,\ldots,1),\\ 2\Delta_{i}\quad&\text{ if }A_{N(i)}=A^{\prime}_{N(i)},\\ 0\quad&\text{ otherwise}.\end{cases}$
We denote $\mathbb{P}_{\mathcal{V}}$ and $\mathbb{P}_{\mathcal{V}^{\prime}}$ the probability measures over the bandit model induced by $T$ rounds of interaction between the policy $\pi$ and the instances $\mathcal{V}$ and $\mathcal{V}^{\prime}$, respectively. In this case, using the divergence decomposition lemma (see Lemma $15.1$ in Lattimore & Szepesvári, 2020) and the fact that the KL-divergence (denoted by $)D_{KL}(.\|.)$ between two $1$-Gaussian distributions with means $\mu_{1}$ and $\mu_{2}$ is equal to $(\mu_{1}-\mu_{2})^{2}/2$, we derive the following equality
$D_{KL}(\mathbb{P}_{\mathcal{V}}\|\mathbb{P}_{\mathcal{V}^{\prime}})=\sum_{i\in[N]}t_{i}(A^{\prime})\frac{(2\Delta_{i})^{2}}{2}=2\sum_{i\in[N]}t_{i}(A^{\prime})\Delta_{i}^{2}.$ (9)
Using Lemma A.1, we have
$D_{KL}(\mathbb{P}_{\mathcal{V}}\|\mathbb{P}_{\mathcal{V}^{\prime}})\leq 2T\left(\sum_{i\in[N]}\frac{\Delta_{i}^{2}}{(k-1)^{d_{i}+1}}\right).$ (10)
For each $t\in[T]$, define the event $\mathcal{E}_{t}$ as:
$\mathcal{E}_{t}\coloneqq\left\{\mathbb{E}\left[\sum_{i\in\mathcal{O}_{t}}\Delta_{i}\right]\geq\frac{1}{2}\sum_{i\in[N]}\Delta_{i}\right\},$
where $\mathcal{O}_{t}\coloneqq\{i\mid A_{tN(i)}=(1,\ldots,1)\}$ denotes the set of units $i$ that receive treatment $1$ at time $t$, along with all their neighbors. This event signifies that in round $t$, the expected sum of $\Delta_{i}$ for such units $i$ is at least half of the total $\sum_{i\in[N]}\Delta_{i}$. We further define the event $\mathcal{E}$ as:
$\mathcal{E}\coloneqq\left\{\sum_{t\in[T]}\mathbbm{1}_{\mathcal{E}_{t}}\geq\frac{T}{2}\right\},$
indicating that $\mathcal{E}_{t}$ occurs in at least half of the $T$ rounds. Using the event $\mathcal{E}$, we can bound the regret of policy $\pi$ for both instances $\mathcal{V}$ and $\mathcal{V}^{\prime}$. For $\mathcal{V}$ we have:
$Reg_{T}(\pi,\mathcal{V})$ $\geq\mathbb{P}_{\mathcal{V}}(\mathcal{E}^{c})\left(Reg_{T}(\pi,\mathcal{V})\mid\mathcal{E}^{c}\right)$
$\geq\mathbb{P}_{\mathcal{V}}(\mathcal{E}^{c})\frac{T}{4N}\sum_{i\in[N]}\Delta_{i},$
where, with a slight abuse of notation, we use $Reg_{T}(\pi,\mathcal{V})\mid\mathcal{E}^{c}$ to denote regret (which is an expectation) conditioned on the event $\mathcal{E}^{c}$. The second line holds because, on $\mathcal{E}^{c}$, the algorithm incurs a regret of at least $\sum_{i\in[N]}\Delta_{i}$ in at least half of the rounds. Consequently, the regret of $\pi$ interacting with $\mathcal{V}$ is at least $\frac{T}{4N}\sum_{i\in[N]}\Delta_{i}$.
Similarly, for $\mathcal{V}^{\prime}$, the regret of $\pi$ can be bounded using $\mathcal{E}$ as:
$Reg_{T}(\pi,\mathcal{V}^{\prime})$ $\geq\mathbb{P}_{\mathcal{V}^{\prime}}(\mathcal{E})\left(Reg_{T}(\pi,\mathcal{V}^{\prime})\mid\mathcal{E}\right)$
$\geq\mathbb{P}_{\mathcal{V}^{\prime}}(\mathcal{E})\frac{T}{4N}\sum_{i\in[N]}\Delta_{i}.$ (11)
On the other hand, Bretagnolle-Huber inequality (Van der Vaart & A. Wellner, 1996) implies:
$\mathbb{P}_{\mathcal{V}}(\mathcal{E}^{c})+\mathbb{P}_{\mathcal{V}^{\prime}}(\mathcal{E})$ $\geq\frac{1}{2}\exp(-D_{KL}(\mathbb{P}_{\mathcal{V}}\|\mathbb{P}_{\mathcal{V}^{\prime}}))$
$\geq\frac{1}{2}\exp{\left(-2T\sum_{i\in[N]}\frac{\Delta_{i}^{2}}{(k-1)^{d_{i}+1}}\right)},$ (12)
where the second line holds by Equation (10). From Equations (A) and (A), we have:
$Reg_{T}(\pi,\mathcal{V})+Reg_{T}(\pi,\mathcal{V}^{\prime})\geq\frac{T}{8N}(\sum_{i\in[N]}\Delta_{i})\exp{\left(-2T\sum_{i\in[N]}\frac{\Delta_{i}^{2}}{(k-1)^{d_{i}+1}}\right)}.$
Next, for each $j\in[M]$ and each $l\in P_{j}$, set the value of $\Delta_{l}$ to:
$\Delta_{l}=\sqrt{\frac{(k-1)^{D_{j}+1}}{TMm_{j}}},$
which satisfies
$T\sum_{i\in[N]}\frac{\Delta_{i}^{2}}{(k-1)^{d_{i}+1}}=T\sum_{j\in[M]}\sum_{l\in P_{j}}\frac{1}{Mm_{j}}=T\sum_{j\in[M]}\frac{1}{M}=1,$
and
$\sum_{i\in[N]}\Delta_{i}$ $=\sqrt{\frac{1}{TM}}\sum_{j\in[M]}\sum_{l\in P_{j}}{\sqrt{\frac{(k-1)^{D_{j}+1}}{m_{j}}}}$
$=\sqrt{\frac{1}{TM}}\sum_{j\in[M]}\sqrt{m_{j}(k-1)^{D_{j}+1}}.$
Note that by the assumption $T\geq\frac{4(k-1)^{\Delta+1}}{M}$, we have $\Delta_{i}\leq\frac{1}{2}$ which implies that the mean rewards are in $[0,1]$.
Now, note that if $k>\frac{2^{\frac{1}{\Delta+1}}}{2^{\frac{1}{\Delta+1}}-1}$, then $(k-1)^{D_{j}+1}>\frac{1}{2}k^{D_{j}+1}$. Using this, there exists a universal constant $C$ such that at least one of $Reg_{T}(\pi,\mathcal{V})$ and $Reg_{T}(\pi,\mathcal{V}^{\prime})$ is greater than
$C\left(\sqrt{\frac{T}{N^{2}M}}\sum_{j\in[M]}\sqrt{m_{j}k^{D_{j}+1}}\right).$
Therefore,
$Reg_{T}(\pi)\in\Omega\left(\sqrt{\frac{T}{N^{2}M}}\sum_{j\in[M]}\sqrt{m_{j}k^{D_{j}+1}}\right).$
[…]
Proof.
Assume $\mathcal{G}^{\prime}$ is obtained by removing an edge between nodes $i$ and $j$. To prove this lemma, consider any instance $\mathcal{V}^{\prime}$ with the interference graph $\mathcal{G}^{\prime}$. Construct an instance $\mathcal{V}$ with the interference graph $\mathcal{G}$ such that the reward distribution for each treatment $A_{N(i)}$ for unit $i$ is identical to its distribution in $\mathcal{V}^{\prime}$.
In this construction, the reward distribution for the unit $i$ becomes independent of the node $j$, and the edge $i-j$ has no impact on the rewards. Consequently, no algorithm can achieve a lower regret on $\mathcal{V}$ compared to $\mathcal{V}^{\prime}$. This implies that for every instance with the interference graph $\mathcal{G}^{\prime}$, a harder instance exists with the interference graph $\mathcal{G}$, which completes the proof. ∎
[…]
Proof.
To prove this lemma, observe that the reward distribution of unit $i$ for each treatment $A_{N(i)}$ assigned to unit $i$ and its neighbors is independent of other treatments. This allows us to construct a corresponding classic multi-armed bandit instance with $k^{d_{i}+1}$ arms, where each arm represents a treatment $A_{N(i)}$ and follows the same reward distribution.
In this scenario, based on the classic lower bound in the multi-armed bandit literature (Lattimore & Szepesvári, 2020), for any policy interacting with a bandit with $K$ arms, there exists an instance where the regret $Reg_{T}(\pi)$ is at least $\sqrt{TK}$. Applying this to our problem, where total regret is defined as the average regret of units, implies that for any policy $\pi$, there exists an instance with the interference graph $\mathcal{G}$ such that:
$Reg_{T}(\pi)=\Omega\left(\sqrt{\frac{Tk^{d_{i}+1}}{N^{2}}}\right).$
[…]
Proof.
For each cluster $C_{i}$, let $|C_{i}|=c_{i}$. It is given that the number of edges between the nodes in $C_{i}$ and $C_{j}$ is at most $r$ for each pair $i,j$. This implies that, within cluster $C_{i}$, there are at most $r(R-1)$ nodes with neighbors outside of $C_{i}$. Let $D_{i}$ denote the set of nodes in $C_{i}$ that have no neighbors outside $C_{i}$. Thus, we have:
$|D_{i}|\geq c_{i}-r(R-1).$
Since $C_{i}$ is a complete graph, it holds that:
$\forall l\in D_{i}:N(l)=C_{i}.$
This indicates that for all nodes $l\in D_{i}$, the set $N(l)$ is identical, which implies that all such nodes belong to the same partition. Furthermore, there are at most $r(R-1)$ nodes in $C_{i}\setminus D_{i}$, meaning that in total these nodes can form:
$\sum_{i\in[R]}|C_{i}\setminus D_{i}|\leq\sum_{i\in[R]}r(R-1)=rR(R-1)$
distinct partitions. This implies that the total number of partitions $M(\mathcal{G})$ is bounded as:
$M(\mathcal{G})\leq R+rR(R-1).$
[…]
Proof.
To prove the lemma, we provide a coloring method using at most $\Delta^{2}+1$ colors, ensuring that two nodes with the same color are neither adjacent nor share a common neighbor. Consider an arbitrary order $\{v_{1},v_{2},\ldots,v_{N}\}$ on the nodes, and let the available colors be $\{c_{1},c_{2},\ldots,c_{\Delta^{2}+1}\}$. Start with $v_{1}$, and for each $v_{i}$, assign the smallest color $c_{j}$ such that no node already colored with $c_{j}$ is adjacent to $v_{i}$ or shares a common neighbor with $v_{i}$.
To prove that this coloring method is valid and does not require more than $\Delta^{2}+1$ colors, assume the process stops at $v_{i}$ because no valid color is available. This would mean $v_{i}$ has more than $\Delta^{2}$ nodes that are either adjacent to it or share a common neighbor. However, since the maximum degree is $\Delta$, $v_{i}$ has at most $\Delta$ neighbors, and each neighbor can have at most $\Delta-1$ other neighbors. This totals at most $\Delta^{2}$ nodes, contradicting the assumption that more than $\Delta^{2}$ nodes are involved. Thus, the coloring method works as intended. This shows that for constant values of $\Delta$ (independent of $N$ and $k$), the value $\Delta^{2}+1$ is also constant.